Solve the following systems of equations for both x and y.
-5x + 3y = -40
9x - 4y = 72

Answer by deborabr(181)

-5x + 3y = -40 *4
9x - 4y = 72 *3
.
-20x + 12y = -160
27x - 12y = 216 suming
__________________
7x +0y = 56
7x = 56
x = 56/7
x = 8
substituting x to find y
-5x + 3y = -40
-5*8 + 3y = -40
-40 + 3y = -40
3y = -40 + 40
3y = 0